# InstructLayout: Instruction-Driven 2D and 3D Layout Synthesis with Semantic Graph Prior

### v1

@@ 4 InstructLayout for 3D Scene Synthesis
Zero-shot Applications Stylization Re-arrangement Completion Uncond.
@@
All experiments are conducted on 3D-FRONT [72], a professionally designed collection of synthetic indoor scenes. However, it does not contain any descriptions of room layouts or object appearances. To construct a high-quality scene-instruction paired dataset, we initially extract view-dependent spatial relations with predefined rules [38; 22]. The dataset is further enhanced by captioning objects with BLIP [27]. To ensure the accuracy of descriptions, the generated captions are filtered by ChatGPT [28; 29] with object ground-truth categories. The final instructions are derived from randomly selected relation triplets. Verbs and conjunctions in the sentences are also randomly picked to keep the diversity and fluency of prompts. Details on 3D dataset curation can be found in Appendix A.1.
@@ 4.3 Instruction-drive 3D Scene Synthesis
Table I presents the quantitive evaluations for synthesizing 3D scenes with instructions. We report the average scores of five runs with different random seeds. As demonstrated, even with the enhancement of quantized semantic features, two baseline methods continue to demonstrate inferior performance compared to ours. ATISS outperforms DiffuScene regarding generation fidelity, owing to its capacity to model in discrete spaces. DiffuScene shows better controllability to ATISS because it affords global visibility of samples during generation. Our proposed InstructLayout exhibits the best of both worlds. Remarkably, we achieve a substantial advancement in controllability, measured in iRecall, for scene generative models, surpassing current state-of-the-art approaches by about 15$\sim$25across various room types, all while maintaining high fidelity. It is noteworthy that InstructLayout excels in handling more complex scenes, such as living and dining rooms, which typically comprise an average of 20 objects, in contrast to bedrooms, which have only 8 objects on average, revealing the benefits of modeling intricate 3D scenes associated with the semantic graph prior. Qualitative visualizations are provided in Appendix B.1.
@@ 4.4 Zero-shot Applications
Stylization and re-arrangement task can be formulated as $p_{\phi}(\mathbf{f}|c,\mathbf{s},\mathbf{y})$ and $p_{\phi,\theta}(\mathbf{s}|c,\mathbf{f},\mathbf{y})$ respectively. In the completion task, we intend to add new objects $\{\mathbf{o}_{k}^{i}\}$ to a partial scene $\mathcal{S}_{i}$ with instructions. By filling the partial scene attributes with [MASK] tokens, we treat them as intermediate states during discrete graph denoising, allowing for a straightforward adaptation of the learned semantic graph prior to these tasks in a zero-shot manner. Unconditional synthesis is implemented by simply setting instruction embeddings $\mathbf{y}$ as zeros. To assess controllability in the stylization task, we define $\Delta\coloneqq\frac{1}{N}\sum_{i=1}^{N}\text{CosSim}(\mathbf{f}_{i},\mathbf{d}^{\text{style}}_{i})-\text{CosSim}(\mathbf{f}_{i},\mathbf{d}^{\text{class}}_{i})$, where $\mathbf{d}^{\text{style}}_{i}$ represents the CLIP text feature of object class name with the desired style, and $\mathbf{d}^{\text{class}}_{i}$ is the CLIP text feature with only class information. $\text{CosSim}(\cdot,\cdot)$ calculates the cosine similarity between two vectors.
Evaluations on zero-shot applications are reported in Table II. Our method consistently outperforms two strong baselines in both controllability and fidelity. While ATISS, as an auto-regressive model, is a natural fit for the completion task, its unidirectional dependency chain limits its effectiveness for tasks requiring global scene modeling, such as re-arrangement. DiffuScene can adapt to these tasks by replacing the known parts with the noised corresponding scene attributes during sampling, similar to image in-painting [77; 78]. However, the known attributes are greatly corrupted in the early steps, which could misguide the denoising direction and necessitate fine-tuning. Additionally, DiffuScene also faces challenges in searching for semantic features in a continuous space for stylization. In contrast, InstructLayout globally models scene attributes and treats partial scene attributes as intermediate discrete states during training. These designs effectively eliminate the training-test gap, rendering it highly versatile for a wide range of downstream tasks. Visualizations of zero-shot applications are available in Appendix B.2.
@@ 5 InstructLayout for 2D Poster Synthesis
AutoPoster [12] is a recently proposed public poster dataset containing 76,537 E-commerce posters with detailed annotations of layout elements. Compared with an ideal comprehensive poster design pipeline, however, it lacks three essential components, including (1) design instructions describing the desired layout of a poster, (2) principle product descriptions for tagline generation, and (3) clean product images for use as poster backgrounds. To the best of our knowledge, no existing dataset offers comprehensive information for the instruction-driven poster synthesis task [41; 10; 43].
In this work, we bridge this gap by constructing a high-quality 2D E-commerce poster dataset based on the AutoPoster dataset. Similar to 3D scene dataset curation, we use predefined rules as a systematic instruction generation pipeline and collect 73,070 product descriptions for AutoPoster posters from public Internet sources using web crawler techniques. For clean product images, previous works[10; 12] employ an off-the-shelf inpainting model LaMa [79] to remove the graphics elements from the poster. However, LaMa often struggles with posters featuring frequent and significant color shifts, resulting in blurry and unnatural distortion of the processed product images. To mitigate this, we first mask out graphic elements with the average color near the masked region and then apply the LaMa inpainting model to obtain clean product images. More details on the 2D poster dataset curation are provided in Appendix A.2.
TABLE III: Quantitive evaluations for instruction-driven 2D poster synthesis on baselines [6; 7; 10; 19] and InstructLayout. Higher iRecall, Val, Undl and Unds, and lower Ove, N-Ali, and Occ indicate better generation quality. The best results among different models are bolded. Standard deviation values are provided as subscripts.
@@ 5.2.1 Baselines
During the re-implementation process, we further condition these baselines with product images and text prompts. While DS-GAN [10] fails to converge with textual instructions, we only feed image conditions to it. Note that most of these methods [6; 7; 10; 57; 19] can only generate categories $c$ and spatial features $\mathbf{s}$, and SAP [12] only predict graphic semantic features $\mathbf{f}$.
@@ 5.2.2 Evaluation Metrics
Besides iRecall proposed in Sec. 4.2.2 to measure the controllability of layout synthesis, we adopt several graphic and content-aware metrics following previous works [10; 42; 41]. Graphic metrics measure the degree of visual harmony among elements (texts, underlay and logos), including Validity (Val) measuring the validity rate of all predicted elements, Overlay (Ove) representing the average Intersection over Union (IoU) for each pair of elements except underlays, Non-Alignment (N-Ali) measuring the extent of spatial non-alignment, and Underlay-Effectiveness (Undl and Unds) indicating the quality of generated underlay. For the content-aware metric, we use Occlusion (Occ) to evaluate the overall visual impact of the poster by measuring the degree of occlusion of the background subject by elements. For semantic feature quality assessment, we follow Lin et al. [12] to calculate the accuracy of predicted dominant colors, stroke colors, and gradient colors in Lab space using Mean Squared Error (MSE) and Mean Absolute Error (MAE).
@@ 5.3 Instruction-drive 2D Poster Synthesis
Table III shows quantitive evaluations on 2D layout generation. We run all models with 5 different random seeds and report the average score. InstructLayout considerably outperforms other existing models by a large margin among all metrics. Especially, our two-stage diffusion model, which skillfully combines discrete and continuous features, remarkably beats both the single-stage diffusion models, the continuous DDPM [57] and the discrete LayoutDM [19], in terms of controllability and visual harmony. It highlights the effectiveness of the proposed discrete graph prior and the capacity of modeling fine spatial relationships facilitated by the continuous diffusion process in the second stage. As presented in Table IV, InstructLayout is also capable of leveraging both image and positional information to mimic the behavior of human designers better and generate graphic semantic features of higher usability. Qualitative results of generated posters with textual content synthesized by the tagline generator are provided in Appendix B.3.
@@ 6.1 Diffusion Timesteps
Although containing two diffusion models, our method could achieve better efficiency by reducing the steps of reverse processes without a noticeable decline in performance. We compare InstructLayout with DiffuScene [21] and LayoutDM [19] in Figure 4 and Table V for 3D scene and 2D poster layout synthesis tasks respectively. We try different timesteps $T$ such as “100+100”, “100+10”, and “25+10”, where the first number represents timesteps for semantic graph prior and the latter is for layout decoder. We find that $T=100$ and $10$ is sufficient for $p_{\phi}(\mathcal{G}|\mathbf{y})$ and $p_{\theta}(\mathcal{S}|\mathcal{G})$ respectively. This stems from the fact that each stage in InstructLayout tackles an easier denoising task compared to single-stage diffusion models such as DiffuScene [21] and LayoutDM [11].
@@ 6.2 Learning Semantic Graph Prior
Permutation Non-invariance. Different to previous studies on graph generation [50; 51; 52], we depart from the permutation-invariant modeling convention to ease the semantic graph prior learning process. We strive to preserve exchangeable graph distributions by randomly shuffling object orders during training. Performance for invariant graph prior is provided in the last column of Table VI. Its performance declines due to the unnecessary imposition of invariance in scene synthesis.
@@ Appendix A Dataset Details
Following previous works [17; 20; 21], we use three types of indoor rooms in 3D-FRONT [72] and preprocess the dataset by filtering some problematic samples, resulting in 4041 bedrooms, 813 living rooms and 900 dining rooms. The number of objects $N_{i}$ in the valid bedrooms is between 3 and 12 with 21 object categories, i.e., $K_{c}=21$. While for living and dining rooms, $N_{i}$ varies from 3 to 21 and $K_{c}=24$. We use the same data split for training and evaluation as ATISS [17].
@@
    Caption Refinement: As generated captions may not always be accurate, we filter them with corresponding ground-truth categories using ChatGPT [28; 29], a large language model fine-tuned for instruction-based tasks. This results in precise and expressive descriptions of each object in the scene. The prompt and hyperparameters for ChatGPT to filter captions are provided in Appendix A.4.
@@
    Instruction Generation: The final instructions for scene synthesis are derived from $1\sim 2$ randomly selected “(subject, relation, object)” triplets obtained during the first extraction process. Verbs and conjunctions within sentences are also randomly picked to maintain diversity and fluency.
@@
    Clean Product Images: In accordance with the improved strategy outlined in Sec. 5.1 for inpainting final posters, we obtained clean product images of significantly higher quality compared to the previous method. The comparisons between previous works [10; 12] are demonstrated in Figure 5.
@@ A.4 Caption Refinement by ChatGPT
Appendix B Qualitative Results
B.1 Instruction-driven 3D Scene Synthesis
We present visualizations of instruction-driven synthesized 3D scenes in Figure 7, 8 and 9. Besides the quantitative evaluations shown in Table I, these qualitative visualizations also show the superiority of our method over previous state-of-the-art approaches in terms of adherence to instructions and generative quality.
@@ B.1 Instruction-driven 3D Scene Synthesis
B.2 Zero-shot 3D Scene Applications
We present visualizations illustrating various zero-shot instruction-driven applications, including stylization, re-arrangement, completion, and unconditional 3D scene synthesis in Figure 10, 11, 12 and 13 respectively. We find that the autoregressive model ATISS [17] tends to generate redundant objects, resulting in chaotic synthesized scenes. DiffuScene [21] encounters challenges in accurately modeling object semantic features, often yielding objects that lack coherence in terms of style or pairing, thereby diminishing the aesthetic appeal of the synthesized scenes. Moreover, both baseline models frequently struggle to follow the provided instructions during conditional generation.
@@ B.2 Zero-shot 3D Scene Applications
B.3 Instruction-driven 2D Poster Synthesis
Comparisons of our synthesized 2D posters with other state-of-the-art methods are shown in Figure 14, 15, and 16, demonstrating the superiority of our method compared to previous state-of-the-art approaches in terms of better mimicking the behavior of human designers and following provided instructions.

### v3

@@ 4 InstructLayout for 3D Scene Synthesis
Zero-shot
Applications
Stylization Re-arrangement Completion Uncond.
@@
All experiments are conducted on 3D-FRONT [71], a professionally designed collection of synthetic indoor scenes. However, it does not contain any descriptions of room layouts or object appearances. To construct a high-quality scene-instruction paired dataset, we initially extract view-dependent spatial relations with predefined rules [37; 21]. The dataset is further enhanced by captioning objects with BLIP [26]. To ensure the accuracy of descriptions, the generated captions are filtered by ChatGPT [27; 28] with object ground-truth categories. The final instructions are derived from randomly selected relation triplets. Verbs and conjunctions in the sentences are also randomly picked to keep the diversity and fluency of prompts. Details on 3D dataset curation can be found in Appendix A.1.
@@ 4.3 Instruction-drive 3D Scene Synthesis
Table I presents the quantitive evaluations for synthesizing 3D scenes with instructions. We report the average scores of five runs with different random seeds. As demonstrated, even with the enhancement of quantized semantic features, two baseline methods continue to demonstrate inferior performance compared to ours. ATISS outperforms DiffuScene regarding generation fidelity, owing to its capacity to model in discrete spaces. DiffuScene shows better controllability to ATISS because it affords global visibility of samples during generation. Our proposed InstructLayout exhibits the best of both worlds. Remarkably, we achieve a substantial advancement in controllability, measured in iRecall, for scene generative models, surpassing current state-of-the-art approaches by about 15$\sim$25across various room types, all while maintaining high fidelity. It is noteworthy that InstructLayout excels in handling more complex scenes, such as living and dining rooms, which typically comprise an average of 20 objects, in contrast to bedrooms, which have only 8 objects on average, revealing the benefits of modeling intricate 3D scenes associated with the semantic graph prior. Qualitative visualizations are provided in Appendix D.1.
@@ 4.4 Zero-shot Applications
[…]
Evaluations on zero-shot applications are reported in Table II. Our method consistently outperforms two strong baselines in both controllability and fidelity. While ATISS, as an auto-regressive model, is a natural fit for the completion task, its unidirectional dependency chain limits its effectiveness for tasks requiring global scene modeling, such as re-arrangement. DiffuScene can adapt to these tasks by replacing the known parts with the noised corresponding scene attributes during sampling, similar to image in-painting [76; 77]. However, the known attributes are greatly corrupted in the early steps, which could misguide the denoising direction and necessitate fine-tuning. Additionally, DiffuScene also faces challenges in searching for semantic features in a continuous space for stylization. In contrast, InstructLayout globally models scene attributes and treats partial scene attributes as intermediate discrete states during training. These designs effectively eliminate the training-test gap, rendering it highly versatile for a wide range of downstream tasks. Visualizations of zero-shot applications are available in Appendix D.2.
@@ 5 InstructLayout for 2D Poster Synthesis
AutoPoster [10] is a recently proposed public poster dataset containing 76,537 E-commerce posters with detailed annotations of layout elements. Compared with an ideal comprehensive poster design pipeline, however, it lacks three essential components, including (1) design instructions describing the desired layout of a poster, (2) principle product descriptions for tagline generation, and (3) clean product images for use as poster backgrounds. To the best of our knowledge, no existing dataset offers comprehensive information for the instruction-driven poster synthesis task [40; 8; 42].
In this work, we bridge this gap by constructing a high-quality 2D E-commerce poster dataset based on the AutoPoster dataset. Similar to 3D scene dataset curation, we use predefined rules as a systematic instruction generation pipeline and collect 73,070 product descriptions for AutoPoster posters from public Internet sources using web crawler techniques. For clean product images, previous works[8; 10] employ an off-the-shelf inpainting model LaMa [78] to remove the graphics elements from the poster. However, LaMa often struggles with posters featuring frequent and significant color shifts, resulting in blurry and unnatural distortion of the processed product images. To mitigate this, we first mask out graphic elements with the average color near the masked region and then apply the LaMa inpainting model to obtain clean product images. More details on the 2D poster dataset curation are provided in Appendix A.2.
TABLE III: Quantitive evaluations for instruction-driven 2D poster synthesis on baselines [4; 5; 8; 18] and InstructLayout. Higher iRecall, Val, Undl and Unds, and lower Ove, N-Ali, and Occ indicate better generation quality. The best results among different models are bolded. Standard deviation values are provided as subscripts.
@@ 5.2.1 Baselines
During the re-implementation process, we further condition these baselines with product images and text prompts. While DS-GAN [8] fails to converge with textual instructions, we only feed image conditions to it. Note that most of these methods [4; 5; 8; 56; 18] can only generate categories $c$ and spatial features $\mathbf{s}$, and SAP [10] only predict graphic semantic features $\mathbf{f}$.
@@ 5.2.2 Evaluation Metrics
Besides iRecall proposed in Sec. 4.2.2 to measure the controllability of layout synthesis, we adopt several graphic and content-aware metrics following previous works [8; 41; 40]. Graphic metrics measure the degree of visual harmony among elements (texts, underlay and logos), including Validity (Val) measuring the validity rate of all predicted elements, Overlay (Ove) representing the average Intersection over Union (IoU) for each pair of elements except underlays, Non-Alignment (N-Ali) measuring the extent of spatial non-alignment, and Underlay-Effectiveness (Undl and Unds) indicating the quality of generated underlay. For the content-aware metric, we use Occlusion (Occ) to evaluate the overall visual impact of the poster by measuring the degree of occlusion of the background subject by elements. For semantic feature quality assessment, we follow Lin et al. [10] to calculate the accuracy of predicted dominant colors, stroke colors, and gradient colors in Lab space using Mean Squared Error (MSE) and Mean Absolute Error (MAE).
@@ 5.3 Instruction-drive 2D Poster Synthesis
Table III shows quantitive evaluations on 2D layout generation. We run all models with 5 different random seeds and report the average score. InstructLayout considerably outperforms other existing models by a large margin among all metrics. Especially, our two-stage diffusion model, which skillfully combines discrete and continuous features, remarkably beats both the single-stage diffusion models, the continuous DDPM [56] and the discrete LayoutDM [18], in terms of controllability and visual harmony. It highlights the effectiveness of the proposed discrete graph prior and the capacity of modeling fine spatial relationships facilitated by the continuous diffusion process in the second stage. As presented in Table IV, InstructLayout is also capable of leveraging both image and positional information to mimic the behavior of human designers better and generate graphic semantic features of higher usability. Qualitative results of generated posters with textual content synthesized by the tagline generator are provided in Appendix D.3.
@@ 6.1 Diffusion Timesteps
Although containing two diffusion models, our method could achieve better efficiency by reducing the steps of reverse processes without a noticeable decline in performance. We compare InstructLayout with DiffuScene [20] and LayoutDM [18] in Figure 4 and Table V for 3D scene and 2D poster layout synthesis tasks respectively. We try different timesteps $T$ such as “100+100”, “100+10”, and “25+10”, where the first number represents timesteps for semantic graph prior and the latter is for layout decoder. We find that $T=100$ and 101010 is sufficient for $p_{\phi}(\mathcal{G}|\mathbf{y})$ and $p_{\theta}(\mathcal{S}|\mathcal{G})$ respectively. This stems from the fact that each stage in InstructLayout tackles an easier denoising task compared to single-stage diffusion models such as DiffuScene [20] and LayoutDM [9].
@@ 6.2 Learning Semantic Graph Prior
Permutation Non-invariance. Different to previous studies on graph generation [49; 50; 51], we depart from the permutation-invariant modeling convention to ease the semantic graph prior learning process. We strive to preserve exchangeable graph distributions by randomly shuffling object orders during training. Performance for invariant graph prior is provided in the last column of Table VI. Its performance declines due to the unnecessary imposition of invariance in scene synthesis.
@@ Appendix A Dataset Details
Following previous works [16; 19; 20], we use three types of indoor rooms in 3D-FRONT [71] and preprocess the dataset by filtering some problematic samples, resulting in 4041 bedrooms, 813 living rooms and 900 dining rooms. The number of objects $N_{i}$ in the valid bedrooms is between 3 and 12 with 21 object categories, i.e., $K_{c}=21$. While for living and dining rooms, $N_{i}$ varies from 3 to 21 and $K_{c}=24$. We use the same data split for training and evaluation as ATISS [16].
@@
    Caption Refinement: As generated captions may not always be accurate, we filter them with corresponding ground-truth categories using ChatGPT [27; 28], a large language model fine-tuned for instruction-based tasks. This results in precise and expressive descriptions of each object in the scene. The prompt and hyperparameters for ChatGPT to filter captions are provided in Appendix A.4.
@@
    Instruction Generation: The final instructions for scene synthesis are derived from 121$\sim$ 21 ∼ 2 randomly selected “(subject, relation, object)” triplets obtained during the first extraction process. Verbs and conjunctions within sentences are also randomly picked to maintain diversity and fluency.
@@
    Clean Product Images: In accordance with the improved strategy outlined in Sec. 5.1 for inpainting final posters, we obtained clean product images of significantly higher quality compared to the previous method. The comparisons between previous works [8; 10] are demonstrated in Figure 6.
@@
A.3 Relation Definations
A.4 Caption Refinement by ChatGPT
Appendix B Computational Costs
Training takes about 2 days on one NVIDIA A40 GPU for the 3D bedroom and 2D poster datasets, and 3 days for the dining room and living room datasets, due to the increased complexity of semantic graphs in larger scenes.
For inference, under the default configuration (diffusion steps $T=100+10$), generating a batch of 128 samples takes approximately 12 seconds for both 2D and 3D layout synthesis. Notably, our method can be significantly accelerated by reducing the number of diffusion steps. For example, setting $T=25+10$ reduces runtime to just 3 seconds, with minimal performance degradation. The impact of the diffusion step count is analyzed in Figure 4 and Table V. We believe that further acceleration is possible through advanced techniques such as flow matching and consistency distillation.
Appendix C User Study
For the evaluation of both 3D scene and 2D poster layout synthesis, a total of 360 questionnaires were gathered from 20 participants. This comprised 200 questionnaires for 3D scene layout synthesis and 160 for 2D poster layout synthesis. In each instance, participants were tasked with selecting the optimal synthesis result based on its fidelity to the given instructions. A method’s User Choice (UC) score is defined as the percentage of times its output was deemed superior by the participants, and the UC scores for all compared methods sum to 100% consequently. As shown in Figure 5, the empirical results demonstrate that InstructLayout achieved markedly higher preference rates over baseline methods, with user preference scores of 87.1% for 3D scene synthesis and 89.5% for 2D poster layout synthesis tasks, respectively.
Figure 5: User study on layout synthesis tasks, encompassing (a) 3D scene generation and (b) 2D poster design.
Appendix D Qualitative Results
D.1 Instruction-driven 3D Scene Synthesis
We present visualizations of instruction-driven synthesized 3D scenes in Figure 8, 9 and 10. Besides the quantitative evaluations shown in Table I, these qualitative visualizations also show the superiority of our method over previous state-of-the-art approaches in terms of adherence to instructions and generative quality.
D.2 Zero-shot 3D Scene Applications
We present visualizations illustrating various zero-shot instruction-driven applications, including stylization, re-arrangement, completion, and unconditional 3D scene synthesis in Figure 11, 12, 13 and 14 respectively. We find that the autoregressive model ATISS [16] tends to generate redundant objects, resulting in chaotic synthesized scenes. DiffuScene [20] encounters challenges in accurately modeling object semantic features, often yielding objects that lack coherence in terms of style or pairing, thereby diminishing the aesthetic appeal of the synthesized scenes. Moreover, both baseline models frequently struggle to follow the provided instructions during conditional generation.
D.3 Instruction-driven 2D Poster Synthesis
Comparisons of our synthesized 2D posters with other state-of-the-art methods are shown in Figure 15, 16, and 17, demonstrating the superiority of our method compared to previous state-of-the-art approaches in terms of better mimicking the behavior of human designers and following provided instructions.
Appendix E Discussions
Out-of-distribution Inputs. For out-of-distribution (OOD) inputs, the proposed system can still produce reasonable layouts, though they may not align closely with the input semantics. For instance, given instructions like “a black cat” or “a football playground”, the generated 3D indoor scenes tend to adopt black tones or football-themed decorations, while still maintaining a plausible bedroom layout, rather than generating a cat or an outdoor playground. Similarly, in 2D poster generation, when provided with natural images, the system interprets them as poster backgrounds and places plausible text bounding boxes accordingly. This generalization ability stems from the pretrained CLIP text and image encoders, which are adapted to extract semantically meaningful features from diverse inputs.
Failure Cases. Although the proposed generative framework demonstrates strong performance and robustness to OOD inputs across both 3D indoor scenes and 2D posters, it still exhibits several failure cases. For example, the predicted 3D bounding boxes do not fully account for the complex relationships between meshes, leading to occasional object collisions in generated 3D scenes, which is also partly due to noisy annotations in the 3D-FRONT dataset. In the 2D poster generation task, the model sometimes produces redundant bounding boxes, especially when the background images already contain text. This can result in visually cluttered designs, as the generated boxes may unintentionally overlap with existing visual elements. Visualization results for these failure cases are provided in Figure 19.
Limitation and Future Work. Another limitation of our work lies in its insufficient generalization across categories for both 3D indoor scene and 2D poster generation. The current approach relies on training a small, specialized generative model for each data type, whereas recent trends favor unified models trained on large-scale, diverse datasets. In light of the rapid development of large vision language models (VLMs), integrating a VLM into our instruction-driven pipeline holds strong potential for improving both generalization and controllability in future work [79; 80]. Meanwhile, designing a unified framework that can generate either a 3D scene [81] or a 2D poster [82] in a single framework is also a promising direction, as it could simplify the pipeline design and accelerate inference.
@@ Appendix E Discussions
(a) Failure cases of 3D scene synthesis: object collisions resulting from limited modeling of inter-object relationships and noisy dataset annotations.
(b) Failure cases of 2D poster synthesis: redundant bounding boxes and overlaps with existing background text due to insufficient content awareness.
Figure 19: Failure cases of InstructLayout.